# Primary gravitational waves at high frequencies I: Origin of suppression in the power spectrum

### v1

@@ Abstract
The primary gravitational waves (PGWs) are generated in the early universe from the quantum vacuum during inflation. In the standard scenario of slow roll inflation, the power spectrum (PS) of PGWs over large scales, which leave the Hubble radius during inflation, is nearly scale-invariant. However, over very small scales, which never leave the Hubble radius, the PS of PGWs behaves as $k^{2}$, where $k$ denotes the wave number. In this work, we examine the PS of PGWs at such high wave numbers or frequencies when the PGWs are evolved post-inflation, through the epochs of radiation and matter domination. Firstly, we argue that the PS has to be regularized in order to truncate the unphysical $k^{2}$ rise at high frequencies. Assuming instantaneous transitions from inflation to the epochs of radiation and matter domination, we carry out the method of adiabatic regularization to arrive at the PS of PGWs over a wide range of frequencies. We show that the process of regularization truncates the $k^{2}$ rise and the PS of PGWs oscillates with a fixed amplitude about a vanishing mean value over small scales or, equivalently, at high frequencies. Secondly, we smooth the transition from inflation to radiation domination (to be precise, we smooth the ‘effective potential’ governing the equation of motion of PGWs) and examine the impact of the smoothing on the regularized PS of PGWs. With the help of a linear smoothing function, we explicitly show that the smoother transition leads to a power-law suppression in the amplitude of the oscillations (about the zero mean value) of the regularized PS of PGWs over small scales that never leave the Hubble radius during inflation. Our analysis indicates that, when transitions are involved, regularization as well as smooth transitions seem essential to ensure that the correlation functions of the PGWs in real space are well behaved. We conclude with a discussion on the different directions in which the results we have obtained need to be extended.
@@ 1 Introduction
where $a_{0}$ is the scale factor of the Friedmann-Lemaître-Robertson-Walker (FLRW) universe today. As convenient, we shall describe the PS of GWs either in terms of wave numbers or frequencies. Note that the CMB constrains the primordial GWs only over large scales corresponding to wave numbers $10^{-4}\lesssim k/a_{0}\lesssim 1\,\rm{Mpc}^{-1}$ or, equivalently, over very small frequencies in the range of $10^{-19}\lesssim f\lesssim 10^{-15}\,\rm{Hz}$. During the past decade, the direct detection of GWs from astrophysical sources, such as coalescing compact binary black holes and neutron stars, by the Ligo-Virgo-Kagra (LVK) collaboration has opened a new window to the universe (in this context, see, for instance, Refs. [32, 33, 34, 35, 36, 37]). The GWs generated in the primordial universe are expected to be stochastic in nature. The detection of GWs from astrophysical sources by the LVK collaboration has led to the expectation that the direct detection of a stochastic GW background (SGWB) may be around the corner. But, despite a few rounds of observations, the LVK collaboration has not yet detected a SGWB over the range of frequencies (say, $1\lesssim f\lesssim 10^{2}\,\mathrm{Hz}$) that its detectors are sensitive to [38, 39, 40]. However, about two years ago, the Pulsar Timing Arrays (PTAs)—such as NANOGrav [41, 42], EPTA (including the data from InPTA) [43, 44], PPTA [45, 46], and CPTA [47]—reported the detection of a SGWB in the range of frequencies $10^{-9}\lesssim f\lesssim 10^{-6}\,\mathrm{Hz}$ [^2] We should mention that the detection of the SGWB by the PTAs also constitute an indirect detection (for a discussion in this context, see, for instance, Ref. [48]).. Many ongoing and forthcoming GW observatories, such as the Square Kilometer Array (SKA) ($10^{-9}\lesssim f\lesssim 10^{-6}\,\mathrm{Hz}$) [49], Laser Interferometer Space Antenna (LISA) ($10^{-4}\lesssim f\lesssim 0.1\,\rm{Hz}$) [50, 51], TianQin and Taiji ($10^{-4}\lesssim f\lesssim 0.1\,\rm{Hz}$) [52, 53, 54], Big Bang Observer (BBO) ($0.1\lesssim f\lesssim 1\,\rm{Hz}$) [55, 56, 57, 58], Decihertz Interferometer Gravitational wave Observatory (DECIGO) ($0.1\lesssim f\lesssim 10\,\rm{Hz}$) [59, 60], Matter-wave Atomic Gradiometer Interferometric Sensor (MAGIS) ($0.1\lesssim f\lesssim 10\,\rm{Hz}$) [61], Cosmic Explorer (CE) ($1\lesssim f\lesssim 10^{2}\,\rm{Hz}$) [62, 63] and Einstein Telescope (ET) ($1\lesssim f\lesssim 10^{3}\,\rm{Hz}$) [64, 65, 66], are expected to operate over the indicated range of frequencies. At high frequencies ($10^{3}\lesssim f\lesssim 10^{20}\,\rm{Hz}$), we already have some upper bounds on the characteristic strain of GWs from Bulk Acoustic Wave devices (BAW) [67, 68], Absolute Radiometer for Cosmology, Astrophysics and Diffuse Emission (ARCADE) [69], Optical Search for QED Vacuum Birefringence, Axions and Photon Regeneration (OSQAR) [70], and the CERN Axion Solar Telescope (CAST) [71]. In addition, a set of detectors have been proposed to observe GWs at these high frequencies (for broad discussions in this context, see, Refs. [72, 73, 74, 75, 76]). Specific detectors include Axion Dark Matter eXperiment (ADMX) [77], Superconducting Quantum Materials and Systems Center (SQMS) [78], International Axion Observatory-heterodyne+single photon detectors (IAXO-HET+SPD) [79], electromagnetic Gaussian beams (GB) [80], Joint Undertaking on the Research for Axion-like particles (JURA) [81], and International Axion Observatory (IAXO) [82]. During the coming decade or two, these observatories are anticipated to provide further constraints on the amplitude and shape of the PS of primordial GWs.
@@
In this work, we shall be interested in understanding the behavior of the PS of PGWs over small scales or, equivalently, at high frequencies. Specifically, we shall be interested in wave numbers that never leave the Hubble radius during the course of cosmological evolution. Let $k_{\mathrm{e}}$ denote the wave number that leaves the Hubble radius at the end of inflation. We shall now briefly discuss the expected behavior of the PS of PGWs over small scales such that $k\gtrsim k_{\mathrm{e}}$. Since these wave numbers remain inside the Hubble radius from the early stages of inflation until today, the mode functions describing the PGWs associated with these wave numbers will always behave in the same manner (barring an overall time-dependent function that depends inversely on the scale factor, say, $a$) as the mode functions in Minkowski spacetime. Then, it can be immediately shown that, over such wave numbers, at any time, the PS of PGWs will behave as $k^{2}$. Moreover, as we shall illustrate in due course, when the PS of PGWs is evaluated today, the $k^{2}$ rise will suggest that their strengths at high frequencies can become comparable to the sensitivities of the ongoing and forthcoming GW observatories (in this context, see Fig. 3). It has been argued that the PS of PGWs have to be regularized, which can truncate the $k^{2}$ rise in these quantities at large wave numbers (see Ref. [83]; for related discussions on the scalar power spectrum, see Refs. [84, 85, 86, 87, 88, 89, 90]; in this context, also see Ref. [91]). As we shall discuss, in the context of PGWs, the process of regularization is essential to avoid the $k^{2}$ rise in the PS on small scales.
@@
Our aim in this work is to carefully examine the effects due to regularization as well as smoothing of the transitions on the PS of PGWs. To understand the behavior around $k_{\mathrm{e}}$, we use the method of adiabatic regularization to calculate the observed PS of PGWs today. We also smooth the transition from inflation to the epoch of radiation domination in a specific manner and explicitly evaluate the corresponding modifications to the PS of PGWs. We show that the regularization leads to a PS which oscillates with a fixed amplitude about a vanishing mean value over $k\gtrsim k_{\mathrm{e}}$. We further illustrate that, at such high wave numbers, the smoother transition leads to a power-law suppression of the amplitude of the oscillations in the regularized PS of PGWs.
This paper is organized as follows. In the following section, we shall describe some of the essential properties of PGWs. We shall describe the equation of motion satisfied by the PGWs, their quantization, and introduce the PS associated with them. We shall also discuss the generic form of the PS expected in the post-inflationary universe. In Sec. 3, we shall discuss the popular case of instantaneous transitions from de Sitter inflation to the epochs of radiation and matter domination, and evaluate the resulting PS of PGWs. In Sec. 4, we shall initially discuss the need for the regularization of the PS of PGWs. We shall also describe the manner in which the PS of PGWs can be regularized using the method of adiabatic subtraction. Moreover, in these sections, we shall clarify a few points related to the amplitude and shape of the actual as well as the regularized PS of PGWs. In particular, we shall show that the regularized PS of PGWs, evaluated during the epochs of radiation and matter domination, oscillates with a constant amplitude about zero over $k\gtrsim k_{\mathrm{e}}$. In Sec. 5, we shall smooth the transition from power-law inflation (in fact, we shall smooth the ‘effective potential’ that governs the dynamics of PGWs) to the epoch of radiation domination with the aid of a linear function and determine the scale factor during the transition. In Sec. 6, we shall evolve the mode functions of the PGWs across the smooth linear transition and illustrate that, over $k\gtrsim k_{\mathrm{e}}$, the regularized PS of PGWs contains oscillations about a vanishing mean and that the amplitude of the oscillations exhibits a power-law suppression. Finally, in Sec. 7, after a quick summary of the results we have obtained, we shall outline a few related issues that need to be examined further.
@@ 2.2 PS of PGWs
where, as is usually done, we have assumed that the expectation value on the left hand side is evaluated in the quantum vacuum. The vacuum state $|0\rangle$ is defined as ${\hat{a}}_{\bm{k}}^{\lambda}|0\rangle=0$ for all $\bm{k}$} and $\lambda$. On utilizing the decomposition (2.4) of the quantized tensor perturbations $\hat{h}_{ij}$ in terms of the rescaled mode functions $\mu_{k}(\eta)$, the PS of PGWs at any time can be expressed as
@@ 2.3 PGWs in the post-inflationary universe
where q<1q<-1 in power-law inflation and $q=-1-\epsilon_{1}$ (with $\epsilon_{1}=-\dot{H}/H^{2}\ll 1$ being the first slow-roll parameter) in slow-roll inflation. Also, $\ell_{0}$ determines the scale of inflation. Moreover, note that exact de Sitter inflation corresponds to q=1q=-1. Further, in the inflationary scenario, it is common to impose the so-called Bunch-Davies initial conditions on the perturbations when the modes are in the sub-Hubble domain during the early stages [100]. This corresponds to assuming that the tensor perturbations are in the vacuum state. Under such assumptions, the solution for the rescaled mode function $\mu_{k}(\eta)$ during inflation can be written as
@@
where $k_{\mathrm{e}}=a_{\mathrm{e}}H_{\mathrm{e}}$, with $a_{\mathrm{e}}$ and $H_{\mathrm{e}}$ denoting the scale factor and the Hubble parameter at the end of inflation. We should emphasize again that the above result is very general and, to arrive at explicit expressions for the PS of PGWs, we need to evaluate the coefficients $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$. In order to calculate these coefficients, we need to specify the manner in which inflation and the epoch of radiation domination are related.
@@ 3.1 PS during radiation domination
We shall now discuss the simplest of scenarios which allows us to calculate the coefficients $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$ easily. We shall consider the popular model wherein de Sitter inflation ends, say, at the conformal time $\eta_{\mathrm{e}}$, and is immediately followed by the epoch of radiation domination. In this scenario, the transition between inflation and the radiation dominated epoch is sharp and reheating is instantaneous. Recall that, during de Sitter inflation and radiation domination, the scale factors are given by [cf. Eqs. (2.18) and (2.12)]
@@
where $a_{\mathrm{eq}}$ denotes the scale factor at the time of radiation-matter equality. Also, $T_{\mathrm{rh}}$ represents the reheating temperature and $H_{0}=2.133\,h\times 10^{-42}\,\mathrm{GeV}$ is the value of the Hubble parameter today. Moreover, $(g_{\ast,\mathrm{rh}},g_{\ast,\mathrm{eq}})$ and $(g_{s,\mathrm{rh}},g_{s,\mathrm{eq}})$ denote the number of effective relativistic degrees of freedom that contribute to the energy density of radiation and entropy at the time of reheating and radiation-matter equality, respectively. The quantities $\Omega_{\mathrm{c}}\,h^{2}=0.12011$, $\Omega_{\mathrm{b}}\,h^{2}=0.0223$, $\Omega_{\gamma}\,h^{2}=2.47\times 10^{-5}$, $\Omega_{\nu}\,h^{2}=1.68\times 10^{-5}$ are the present day values for the dimensionless densities of cold dark matter, baryons, photons and neutrinos, respectively [101]. Further, to evaluate the argument of the trigonometric functions appearing in the PS, we need to evaluate the quantity $x$. Recall that

### v2

@@ Abstract
The primary gravitational waves (PGWs) are generated in the early universe from the quantum vacuum during inflation. In the standard scenario of slow roll inflation, the power spectrum (PS) of PGWs over large scales, which leave the Hubble radius during inflation, is nearly scale-invariant. However, over very small scales, which never leave the Hubble radius, the PS of PGWs behaves as $k^{2}$, where $k$ denotes the wave number. In this work, we examine the PS of PGWs at such high wave numbers or frequencies when the PGWs are evolved post-inflation, through the epochs of radiation and matter domination. Firstly, we argue that the PS has to be regularized in order to truncate the unphysical $k^{2}$ rise at high frequencies. Assuming instantaneous transitions from inflation to the epochs of radiation and matter domination, we carry out the method of adiabatic regularization to arrive at the PS of PGWs over a wide range of frequencies. We show that the process of regularization truncates the $k^{2}$ rise and that the regularized PS of PGWs oscillates with a fixed amplitude about a vanishing mean value over small scales or, equivalently, at high frequencies. Secondly, we smooth the transition from inflation to radiation domination (to be precise, we smooth the ‘effective potential’ governing the equation of motion of PGWs) and examine the impact of the smoothing on the regularized PS of PGWs. With the help of a linear smoothing function, we explicitly show that the smoother transition leads to a power-law suppression in the amplitude of the oscillations (about the zero mean value) of the regularized PS of PGWs over small scales that never leave the Hubble radius during inflation. Our analysis indicates that, when transitions are involved, regularization as well as smooth transitions seem essential to ensure that the correlation functions of the PGWs in real space are well behaved. We conclude with a discussion on the different directions in which the results we have obtained need to be extended.
@@ 1 Introduction
where $a_{0}$ is the scale factor of the Friedmann-Lemaître-Robertson-Walker (FLRW) universe today. As convenient, we shall describe the PS of GWs either in terms of wave numbers or frequencies. Note that the CMB constrains the primordial GWs only over large scales corresponding to wave numbers $10^{-4}\lesssim k/a_{0}\lesssim 1\,\rm{Mpc}^{-1}$ or, equivalently, over very small frequencies in the range of $10^{-19}\lesssim f\lesssim 10^{-15}\,\rm{Hz}$. During the past decade, the direct detection of GWs from astrophysical sources, such as coalescing compact binary black holes and neutron stars, by the Ligo-Virgo-Kagra (LVK) collaboration has opened a new window to the universe (in this context, see, for instance, Refs. [4, 3, 8, 7, 5, 2]). The GWs generated in the primordial universe are expected to be stochastic in nature. The detection of GWs from astrophysical sources by the LVK collaboration has led to the expectation that the direct detection of a stochastic GW background (SGWB) may be around the corner. But, despite a few rounds of observations, the LVK collaboration has not yet detected a SGWB over the range of frequencies (say, $1\lesssim f\lesssim 10^{2}\,\mathrm{Hz}$) that its detectors are sensitive to [9, 6, 10]. However, about two years ago, the Pulsar Timing Arrays (PTAs)—such as NANOGrav [17, 18], EPTA (including the data from InPTA) [26, 27], PPTA [138, 115], and CPTA [135]—reported the detection of a SGWB in the range of frequencies $10^{-9}\lesssim f\lesssim 10^{-6}\,\mathrm{Hz}$ [^2] We should mention that the detection of the SGWB by the PTAs also constitute an indirect detection (for a discussion in this context, see, for instance, Ref. [136]).. Many ongoing and forthcoming GW observatories, such as the Square Kilometer Array (SKA) ($10^{-9}\lesssim f\lesssim 10^{-6}\,\mathrm{Hz}$) [86], Laser Interferometer Space Antenna (LISA) ($10^{-4}\lesssim f\lesssim 0.1\,\rm{Hz}$) [31, 28], TianQin and Taiji ($10^{-4}\lesssim f\lesssim 0.1\,\rm{Hz}$) [84, 101, 76], Big Bang Observer (BBO) ($0.1\lesssim f\lesssim 1\,\rm{Hz}$) [54, 53, 82, 29], Decihertz Interferometer Gravitational wave Observatory (DECIGO) ($0.1\lesssim f\lesssim 10\,\rm{Hz}$) [91, 90], Matter-wave Atomic Gradiometer Interferometric Sensor (MAGIS) ($0.1\lesssim f\lesssim 10\,\rm{Hz}$) [52], Cosmic Explorer (CE) ($1\lesssim f\lesssim 10^{2}\,\rm{Hz}$) [67, 68] and Einstein Telescope (ET) ($1\lesssim f\lesssim 10^{3}\,\rm{Hz}$) [121, 41, 1], are expected to operate over the indicated range of frequencies. At high frequencies ($10^{3}\lesssim f\lesssim 10^{20}\,\rm{Hz}$), we already have some upper bounds on the characteristic strain of GWs from Bulk Acoustic Wave devices (BAW) [77, 74], Absolute Radiometer for Cosmology, Astrophysics and Diffuse Emission (ARCADE) [70], Optical Search for QED Vacuum Birefringence, Axions and Photon Regeneration (OSQAR) [30], and the CERN Axion Solar Telescope (CAST) [139]. In addition, a set of detectors have been proposed to observe GWs at these high frequencies (for broad discussions in this context, see, Refs. [127, 62, 42, 87, 89]). Specific detectors include Axion Dark Matter eXperiment (ADMX) [32], Superconducting Quantum Materials and Systems Center (SQMS) [37], International Axion Observatory-heterodyne+single photon detectors (IAXO-HET+SPD) [116], electromagnetic Gaussian beams (GB) [94], Joint Undertaking on the Research for Axion-like particles (JURA) [36], and International Axion Observatory (IAXO) [120]. During the coming decade or two, these observatories are anticipated to provide further constraints on the amplitude and shape of the PS of primordial GWs (in this context, also see Ref. [93]).
@@
In this work, we shall be interested in understanding the behavior of the PS of PGWs over small scales or, equivalently, at high frequencies. Specifically, we shall be interested in wave numbers that never leave the Hubble radius during the course of cosmological evolution. Let $k_{\mathrm{e}}$ denote the wave number that leaves the Hubble radius at the end of inflation. We shall now briefly discuss the expected behavior of the PS of PGWs over small scales such that $k\gtrsim k_{\mathrm{e}}$. Since these wave numbers remain inside the Hubble radius from the early stages of inflation until today, the Fourier mode functions describing the PGWs associated with these wave numbers will always behave in the same manner (barring an overall time-dependent function that depends inversely on the scale factor, say, $a$) as the mode functions in Minkowski spacetime. Then, it can be immediately shown that, over such wave numbers, at any time, the PS of PGWs will behave as $k^{2}$. Moreover, as we shall illustrate in due course, when the PS of PGWs is evaluated today, the $k^{2}$ rise will suggest that their strengths at high frequencies can become comparable to the sensitivities of the ongoing and forthcoming GW observatories (in this context, see Fig. 3). It has been argued that the PS of PGWs has to be regularized, which can truncate the $k^{2}$ rise in the quantity at large wave numbers (see Ref. [131]; for related discussions on the scalar power spectrum, see Refs. [112, 20, 21, 22, 128, 58, 114]; in this context, also see Ref. [69]). As we shall discuss, in the context of PGWs, the process of regularization is essential to avoid the $k^{2}$ rise in the PS on small scales.
@@
Our aim in this work is to carefully examine the effects due to regularization as well as smoothing of the transitions on the PS of PGWs. To understand the behavior around $k_{\mathrm{e}}$, we use the method of adiabatic regularization to calculate the observed PS of PGWs today. We also smooth the transition from inflation to the epoch of radiation domination in a specific manner and explicitly evaluate the corresponding modifications to the PS of PGWs. We show that, in the case of an instantaneous transition, regularization leads to a PS which oscillates with a fixed amplitude about a vanishing mean value over $k\gtrsim k_{\mathrm{e}}$. We further illustrate that, at such high wave numbers, the smoother transition leads to a power-law suppression of the amplitude of the oscillations in the regularized PS of PGWs.
This paper is organized as follows. In the following section, we shall describe some of the essential properties of PGWs. We shall describe the equation of motion satisfied by the PGWs, their quantization, and introduce the PS associated with them. We shall also discuss the generic form of the PS expected in the post-inflationary universe. In Sec. 3, we shall discuss the popular case of instantaneous transitions from de Sitter inflation to the epochs of radiation and matter domination, and evaluate the resulting PS of PGWs. In Sec. 4, we shall initially discuss the need for the regularization of the PS of PGWs. We shall also describe the manner in which the PS of PGWs can be regularized using the method of adiabatic subtraction. Moreover, in these sections, we shall clarify a few points related to the amplitude and shape of the actual as well as the regularized PS of PGWs. In particular, we shall show that, the regularized PS of PGWs, evaluated during the epochs of radiation and matter domination, oscillates with a constant amplitude about zero over $k\gtrsim k_{\mathrm{e}}$. In Sec. 5, we shall smooth the transition from power-law inflation (in fact, we shall smooth the ‘effective potential’ that governs the dynamics of PGWs) to the epoch of radiation domination with the aid of a linear function and determine the scale factor during the transition. In Sec. 6, we shall evolve the mode functions of the PGWs across the smooth linear transition and illustrate that, during radiation domination, over $k\gtrsim k_{\mathrm{e}}$, the regularized PS of PGWs contains oscillations about a vanishing mean and that the amplitude of the oscillations exhibits a power-law suppression. Finally, in Sec. 7, after a quick summary of the results we have obtained, we shall outline a few related issues that need to be examined further.
@@ 2.2 PS of PGWs
where, as is usually done, we have assumed that the expectation value on the left hand side is evaluated in the quantum vacuum $|0\rangle$, which is defined as ${\hat{a}}_{\bm{k}}^{\lambda}|0\rangle=0$ for all $\bm{k}$} and $\lambda$. On utilizing the decomposition (2.4) of the quantized tensor perturbations $\hat{h}_{ij}$ in terms of the rescaled mode functions $\mu_{k}(\eta)$, the PS of PGWs at any time can be expressed as
@@ 2.3 PGWs in the post-inflationary universe
where q<1q<-1 in power-law inflation and $q=-1-\epsilon_{1}$ (with $\epsilon_{1}=-\dot{H}/H^{2}\ll 1$ being the first slow roll parameter) in slow roll inflation. Also, $\ell_{0}$ determines the scale of inflation. Moreover, note that exact de Sitter inflation corresponds to q=1q=-1. Further, in the inflationary scenario, it is common to impose the so-called Bunch-Davies initial conditions on the perturbations when the modes are in the sub-Hubble domain during the early stages [44, 46, 45]. This corresponds to assuming that the tensor perturbations are in the vacuum state. Under such assumptions, the solution for the rescaled mode function $\mu_{k}(\eta)$ during inflation can be written as
@@
where $k_{\mathrm{e}}=a_{\mathrm{e}}H_{\mathrm{e}}$, with $a_{\mathrm{e}}$ and $H_{\mathrm{e}}$ denoting the scale factor and the Hubble parameter at the end of inflation. We should emphasize again that the above result is very general and, to arrive at explicit expressions for the PS of PGWs, we need to evaluate the coefficients $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$. In order to calculate these coefficients, we need to specify the manner in which inflation and the epoch of radiation domination are related. There is a point we wish to make before we proceed further. Note that, if we assume that inflation is of the de Sitter form (so that q=1q=-1), then the rescaled mode function [cf. Eqs. (2.19) and (2.20)] simplifies to be
$\mu_{k}(\eta)=\frac{1}{\sqrt{2k}}\left(1-\frac{i}{k\eta}\right)\mathrm{e}^{-ik\eta}.$ (2.25)
In such a case, at the end of inflation, say, at the conformal time $\eta_{\mathrm{e}}$, the PS of PGWs can be expressed as
$\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta_{\mathrm{e}})=\frac{2H_{\mathrm{e}}^{2}}{\pi^{2}M_{{}_{\mathrm{Pl}}}^{2}}\left(1+y_{\mathrm{e}}^{2}\right),$ (2.26)
where $y_{\mathrm{e}}=-k\eta_{\mathrm{e}}=k/k_{\mathrm{e}}$. This spectrum is scale invariant over $y_{\mathrm{e}}\ll 1$, whereas it behaves as $y_{\mathrm{e}}^{2}$ over $y_{\mathrm{e}}\gg 1$. We shall see that, over $y_{\mathrm{e}}\gg 1$, the PS continues to behave as $y_{\mathrm{e}}^{2}$ even when we evolve the modes post-inflation.
@@ 3.1 PS during radiation domination
We shall now discuss the simplest of scenarios which allows us to calculate the coefficients $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$ easily. We shall consider the popular model wherein de Sitter inflation ends at the conformal time $\eta_{\mathrm{e}}$, and is immediately followed by the epoch of radiation domination. In this scenario, the transition between inflation and the radiation-dominated epoch is sharp and reheating is instantaneous. Recall that, during de Sitter inflation and radiation domination, the scale factors are given by [cf. Eqs. (2.18) and (2.12)]
@@
where $a_{\mathrm{eq}}$ denotes the scale factor at the time of radiation-matter equality. Also, $T_{\mathrm{rh}}$ represents the reheating temperature and $H_{0}=2.133\,h\times 10^{-42}\,\mathrm{GeV}$ is the value of the Hubble parameter today. Moreover, $(g_{\ast,\mathrm{rh}},g_{\ast,\mathrm{eq}})$ and $(g_{s,\mathrm{rh}},g_{s,\mathrm{eq}})$ denote the number of effective relativistic degrees of freedom that contribute to the energy density and entropy of radiation at the time of reheating and radiation-matter equality, respectively. The quantities $\Omega_{\mathrm{c}}\,h^{2}=0.12011$, $\Omega_{\mathrm{b}}\,h^{2}=0.0223$, $\Omega_{\gamma}\,h^{2}=2.47\times 10^{-5}$, $\Omega_{\nu}\,h^{2}=1.68\times 10^{-5}$ are the present day values for the dimensionless densities of cold dark matter, baryons, photons and neutrinos, respectively [19]. Further, to evaluate the argument of the trigonometric functions appearing in the PS, we need to calculate the quantity $x$. Recall that